### v2

@@ 5.1 Point of origin
NGC 4522 was mapped in VIVA and is a well-known example of a galaxy undergoing ram pressure stripping (e.g. Kenney & Koopmann, 1999; Kenney et al., 2004; Vollmer et al., 2004; Chung et al., 2009). Furthermore, the direction of the stripping appears to be aligned along the direction towards BC6. However, BC6 and NGC 4522 are separated by 1800 km/s along the line-of-sight. Vollmer et al. (2006) attempted to simulate the gas stripping occurring in NGC 4522, finding that it was best fit by a very high velocity relative to the ICM (3500 km s), with a velocity component in the SE direction of 3000 km s. At this velocity the distance between NGC 4522 and BC6 could be traversed in 100 Myr. Vollmer et al. (2006) estimate that the period of maximum stripping pressure on NGC 4522 was approximately 50 Myr ago. These two time frames do not exactly line up, but Vollmer et al. (2006) indicated that significant stripping is expected to begin well before the peak ram pressure is reached. With these high velocities it may also be possible to explain the large velocity offset between BC6 and NGC 4522, if there is a significant component of NGC 4522’s velocity (relative to the ICM) along the line-of-sight.
@@
The second issue raised by Minchin et al. (2019) is that it is unlikely that the local ICM radial velocity in this region of Virgo is as low as 500 km s (recall that in §2 we mentioned that the typical velocity of galaxies in this regions of Virgo is 1000 km s). If this is not the case then it is not plausible that the H i gas could have initially had a radial velocity of 2300 km s and then slowed to 500 km s. This is a more serious objection to NGC 4522 as the progenitor and is difficult to dismiss. However, we note that a similar (though less severe) objection applies to all the candidate progenitor galaxies to the NW and SE listed in Table 3, as these all have radial velocities above 500 km s. If one of these is indeed the origin of BC6, then this may be an indication that we do not understand the motions (and perhaps, sub-structure) of the ICM in this region of Virgo.
@@
To summarize, we have considered a number of potential progenitor galaxies within a few degrees of BC6, but there is no clear parent object, in fact, all the candidates appear to have at least one disqualifying factor. We consider NGC 4522 to be the most likely parent object because it has similar metallicity and is a known case of extreme ram pressure stripping with the gas tail pointing towards BC6. However, it is difficult to account for the kinematic separation of NGC 4522 and BC6. NGC 4380 and NGC 4519 also remain candidates for the progenitor, but this may simply be because there are fewer observations of these galaxies and we could be ignorant of disqualifying information. NGC 4445 was suggested as the point of origin by Kent et al. (2009) and Minchin et al. (2019), but we disfavor this as Sorgho et al. (2017) seem to show that it has a gas tail that points away from BC6 and because the observed H i complex represents almost the entirety of its estimated original gas content.
@@ 5.2 Fate
and is the temperature of the hot ICM, is its number density (in cm), and is the cloud radius (in pc). To evaluate this and proceed we need to know the physical properties of both a cloud within the H i complex and the surrounding ICM.
To obtain approximate physical properties of a single cloud (as far as possible) we take the values for C2 (Figure 3, left) which was observed with the VLA in Kent et al. (2009). In this higher resolution imaging the cloud has an elongated structure with two main concentrations of mass. For simplicity we take a quarter of the H i major axis (1.5′) measured as the radius of a single cloud (2 kpc) and half of the total H i mass of C2 as its mass ( M). Here we have also multiplied by 1.4 to account for Helium and other metals. Assuming a spherical geometry gives a cloud number density of 0.3 .
For the ICM properties we use the simple -model ICM density profile used in Vollmer et al. (2001), evaluated at 2.7 (the separation between BC6 and M 87), to obtain a ICM density estimate of cm. For the temperature we see from the profile plotted by Böhringer et al. (1994) that at 2.7 from M 87 the average temperature of the ICM is roughly 2 keV ( K).
@@
where is the average number density of the cloud (in cm). The other quantity are the same as in Equation 1. Evaluating this expression gives a cloud lifetime for C2 of over a Gyr.
Although this is only an order of magnitude estimate it is clear from the stellar population (§4.1) that the H i complex has been traversing the ICM for at least 100 Myr and is still an extremely rich H i structure. Furthermore, Calura et al. (2020) considered in detail the survival of a much smaller gas cloud M (designed to mimic SECCO 1), finding that when simulating it traversing the ICM at high speed it could still survive as a distinct structure for around a Gyr. Thus, given the enormous reservoir still present, it seems that BC6 is not in imminent danger of losing its H i gas and is unlikely to become unbound and disperse into the cluster in the short term.
@@ 6 Conclusions
We have identified the stellar counterpart (BC6) of an enormous, isolated H i structure (190 kpc long, M) in the Virgo cluster that has been thought to be “dark” since its discovery by Kent et al. (2007). The extremely low (stellar) mass counterpart lies at the NW tip of the structure (Figure 3). In ground-based imaging it appears as faint and irregular blue clumps, which HST images reveal to consist exclusively of young (100 Myr) stars. Given this population, the total stellar mass of BC6 is likely only a few times M, making it a truly extreme object in terms of gas fraction.
An optical spectrum of this counterpart from HET returned an H redshift (500 km s) consistent with the velocity of the co-spatial H i gas seen in ALFALFA and with GBT observations, thereby confirming association. This spectrum also revealed that BC6 is metal-rich (given its low stellar mass), indicating that its gas complex must have originated in a galaxy where the gas could be pre-enriched.
@@ Appendix A Other counterpart candidates
Although our citizen science search for additional blue blob candidates throughout Virgo is incomplete at the time of writing, we extracted all the candidates that currently exist in this region to identify if BC6 is the only object associated with the H i tail, or if there might be other candidate blue blobs coincident with other clouds.
We only found one additional candidate (Figure 6) coincident with the H i structure, next to cloud C4 (Figure 3). The appearance of this candidate suggests that it is likely a background group of interacting galaxies. Although there are blue clumps, as in genuine blue blobs, they are surrounded by diffuse yellow light, probably made up of old stars that are very far away. Thus, we do not consider this likely to be a genuine blob blue or associated with the H i structure, however, we plan to obtain optical spectroscopic follow-up in the future to rule this out.

### v3

@@ 5.1 Point of origin
NGC 4522 was mapped in VIVA and is a well-known example of a galaxy undergoing ram pressure stripping (e.g. Kenney & Koopmann, 1999; Kenney et al., 2004; Vollmer et al., 2004; Chung et al., 2009). Furthermore, the direction of the stripping appears to be aligned along the direction towards BC6. However, BC6 and NGC 4522 are separated by 1800 km/s along the line-of-sight. Vollmer et al. (2006) attempted to simulate the gas stripping occurring in NGC 4522, finding that it was best fit by a very high velocity relative to the ICM (3500 km s-1), with a velocity component in the SE direction of 3000 km s-1. At this velocity the distance between NGC 4522 and BC6 could be traversed in 100 Myr. Vollmer et al. (2006) estimate that the period of maximum stripping pressure on NGC 4522 was approximately 50 Myr ago. These two time frames do not exactly line up, but Vollmer et al. (2006) indicated that significant stripping is expected to begin well before the peak ram pressure is reached. With these high velocities it may also be possible to explain the large velocity offset between BC6 and NGC 4522, if there is a significant component of NGC 4522’s velocity (relative to the ICM) along the line-of-sight. Indeed, Minchin et al. (2019) identified a low column density extension of NGC 4522’s ram pressure tail pointing in the direction of BC6 and the H i complex. Although, this makes it tempting to conclude that the two are likely connected, this only extends about 5′ and 200 km s-1 towards the complex and therefore cannot be considered conclusive. Having said this, extrapolating this same velocity gradient another 40′ would bring us to both the approximate position and velocity of the H i complex.
@@
The second issue raised by Minchin et al. (2019) is that it is unlikely that the local ICM radial velocity in this region of Virgo is as low as 500 km s-1 (recall that in §2 we mentioned that the typical velocity of galaxies in this regions of Virgo is 1000 km s-1). If this is not the case then it is not plausible that the H i gas could have initially had a radial velocity of 2300 km s-1 and then slowed to 500 km s-1. This is a more serious objection to NGC 4522 as the progenitor and is difficult to dismiss. The ram pressure stripping simulations of Vollmer et al. (2006) are best able to reproduce the extraplanar gas distribution of NGC 4522 with an incidence angle of about 60∘ between the galaxy’s motion and the ICM’s. However, in this configuration the radial velocity of the ICM would be approximately 1000 km s-1. An ICM velocity as low as 500 km s-1 could be possible in different configurations, but Vollmer et al. (2006) found that such configurations would not reproduce the observed gas distribution. We note that a similar (though less severe) objection regarding the ICM velocity applies to all the candidate progenitor galaxies to the NW and SE listed in Table 3, as these all have radial velocities above 500 km s-1. If one of these is indeed the origin of BC6, then this may be an indication that we do not understand the motions (and perhaps, sub-structure) of the ICM in this region of Virgo, and it might be necessary to re-visit the hydrodynamical modeling of the gas stripping in NGC 4522 to determine if other factors such as irregular motions and clumpiness in the ICM could reconcile these differences.
@@
To summarize, we have considered a number of potential progenitor galaxies within a few degrees of BC6, but there is no clear parent object, in fact, all the candidates appear to have at least one disqualifying factor. We consider NGC 4522 to be the most likely parent object because it has similar metallicity and is a known case of extreme ram pressure stripping with the gas tail pointing towards BC6 both in projection and velocity. However, it is difficult to account for the kinematic separation of NGC 4522 and BC6. NGC 4380 and NGC 4519 also remain candidates for the progenitor, but this may simply be because there are fewer observations of these galaxies and we could be ignorant of disqualifying information. NGC 4445 was suggested as the point of origin by Kent et al. (2009) and Minchin et al. (2019), however, we slightly disfavor this galaxy as the progenitor as the observed H i complex represents almost the entirety of its estimated original gas content and the ram pressure tail would need a particularly wide opening angle to explain the extent of the complex.
@@ 5.2 Fate
and is the temperature of the hot ICM, is its number density (in cm-3), and is the cloud radius (in pc). To evaluate this and proceed we need to know the physical properties of both a cloud within the H i complex and the surrounding ICM.
To obtain approximate physical properties of a single cloud (as far as possible) we take the values for C2 (Figure 3, left) which was observed with the VLA in Kent et al. (2009). In this higher resolution imaging the cloud has an elongated structure with two main concentrations of mass. For simplicity we take a quarter of the H i major axis (1.5′) measured as the radius of a single cloud (2 kpc) and half of the total H i mass of C2 as its mass ( M⊙). Here we have also multiplied by 1.4 to account for Helium and other metals. Assuming a spherical geometry gives a cloud number density of 0.3 .
For the ICM properties we use the simple -model ICM density profile used in Vollmer et al. (2001), evaluated at 2.7∘ (the separation between BC6 and M 87), to obtain a ICM density estimate of cm-3. For the temperature we see from the profile plotted by Böhringer et al. (1994) that at 2.7∘ from M 87 the average temperature of the ICM is roughly 2 keV ( K).
@@
where is the average number density of the cloud (in cm-3). The other quantity are the same as in Equation 1. Evaluating this expression gives a cloud lifetime for C2 of over a Gyr.
Although this is only an order of magnitude estimate it is clear from the stellar population (§4.1) that the H i complex has been traversing the ICM for at least 100 Myr and is still an extremely rich H i structure. Furthermore, Calura et al. (2020) considered in detail the survival of a much smaller gas cloud M⊙ (designed to mimic SECCO 1), finding that when simulating it traversing the ICM at high speed it could still survive as a distinct structure for around a Gyr. Thus, given the enormous reservoir still present, it seems that BC6 is not in imminent danger of losing its H i gas and is unlikely to become unbound and disperse into the cluster in the short term.
@@ 6 Conclusions
We have identified the stellar counterpart (BC6) of an enormous, isolated H i structure (190 kpc long, M⊙) in the Virgo cluster that has been thought to be “dark” since its discovery by Kent et al. (2007). The extremely low (stellar) mass counterpart lies at the NW tip of the structure (Figure 3). In ground-based imaging it appears as faint and irregular blue clumps, which HST images reveal to consist exclusively of young (100 Myr) stars. Given this population, the total stellar mass of BC6 is likely only a few times M⊙, making it a truly extreme object in terms of gas fraction.
An optical spectrum of this counterpart from HET returned an H redshift (500 km s-1) consistent with the velocity of the co-spatial H i gas seen in ALFALFA and with GBT observations, thereby confirming association. This spectrum also revealed that BC6 is metal-rich (given its low stellar mass), indicating that its gas complex must have originated in a galaxy where the gas could be pre-enriched.
@@ Appendix A Other counterpart candidates
Our citizen science search for additional blue blob candidates throughout Virgo was recently completed, we extracted all the candidates in this region to identify if BC6 is the only object associated with the H i tail, or if there might be other candidate blue blobs coincident with other clouds.
We only found one additional candidate (Figure 7) coincident with the H i structure, next to cloud C4 (Figure 3). The appearance of this candidate suggests that it is likely a background group of interacting galaxies. Although there are blue clumps, as in genuine blue blobs, they are surrounded by diffuse yellow light, probably made up of old stars that are very far away. Thus, we do not consider this likely to be a genuine blue blob or associated with the H i structure, however, we plan to obtain optical spectroscopic follow-up in the future to rule this out.